hi can you please help me solve linear equation i cant understant them like 5x+2y=10 and x-2y=2
Answer by Alan3354(69443)

5x+2y=10
x -2y =2
-------------- Add
6x = 12
x = 2
-----
x -2y =2
2 -2y =2
y = 0